OpenStudy (anonymous):

how do i turn 6.004 into a percent?

OpenStudy (anonymous):

Divide it but 100.

OpenStudy (anonymous):

by* not but

myininaya (myininaya):

lol

OpenStudy (anonymous):

lol so .6745 just divide by 100

myininaya (myininaya):

1=100% so if you want to change something to percent multiply it by 100%

myininaya (myininaya):

like .5=50%

myininaya (myininaya):

.75=75%

myininaya (myininaya):

all of these numbers are being multiplied by 100%

myininaya (myininaya):

s0 6.004=6.004*1=6.004*100%=600.4%

myininaya (myininaya):

you are actually just moving the decimal over twice

myininaya (myininaya):

do you see?

OpenStudy (anonymous):

i see so what bout .6745

myininaya (myininaya):

i gave you 4 examples see if you can get right by urself. i will say if it is right or wrong

OpenStudy (anonymous):

674.5/100

myininaya (myininaya):

ok .75=75% we should moved the decimal over twice (we are not dividing we are multiplying)

myininaya (myininaya):

ignore what you wrote we have .6745 just move the decimal over twice and put a percent sign on the number

OpenStudy (anonymous):

Are you trying to find the percentage of a certain number or are you just trying to convert the numbers to percentages? If you're trying to convert to percentages just move the decimal to the right 2 places and put the % sign on it.

myininaya (myininaya):

.5252352=52.52352% .64=64% .642=64.2%

OpenStudy (anonymous):

Are you trying to find the percentage of a certain number or are you just trying to convert the numbers to percentages? If you're trying to convert to percentages just move the decimal to the right 2 places and put the % sign on it.

myininaya (myininaya):

.8241=82.41% .99912=99.912%

myininaya (myininaya):

do you see a pattern?

OpenStudy (anonymous):

sorry got tied up

OpenStudy (anonymous):

what about writing percents into fractions

OpenStudy (anonymous):

put them over 100.

OpenStudy (anonymous):

just makin sure thank u